OpenStudy (anonymous):

the square root of 36/49

OpenStudy (anonymous):

6/7

Parth (parthkohli):

\[\sqrt{36 \over 49} \implies {\sqrt{36} \over \sqrt{49}}\]

OpenStudy (anonymous):

\[\sqrt{\frac{a}{b}} = \frac{\sqrt{a}}{\sqrt{b}}\]

OpenStudy (anonymous):

all of you :)